I am completely in the dark with application problems,,,,,,
here is one,,,,
the power rating P in watts, of an electrical appliance is determined bu p=IxV, where I is the cuurent in amperes and V is measured in volts.
A. A kitchen requires 30 amps of current and the coltage in the house is 115 volts, wht is the wattage of the kitchen?
b. Solve for I;for V.
need answer by tomorrow ,,,thanks

Answer by mananth(16946)


the power rating P in watts, of an electrical appliance is determined bu p=IxV, where I is the cuurent in amperes and V is measured in volts.
A. A kitchen requires 30 amps of current and the coltage in the house is 115 volts, wht is the wattage of the kitchen?
b. Solve for I;for V.

p = I*V
P= 30 * 115 = 345 watts